A potential difference of 12 V produces a current of 0.40 A in a piece of copper wire. What is the resistance of the wire ?
1
Expert's answer
2012-09-14T12:07:27-0400
Let's use the Ohm's law:

I = U/R ==> R = U/I = 12V/0.4A = 30 Ω.